OpenStudy (anonymous):

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions. 3x^2+6x-5=0

OpenStudy (bahrom7893):

D = b^2-4ac = 36 - 4*3*(-5) = 36 + 60 = 96

OpenStudy (bahrom7893):

Two different irrational solutions

OpenStudy (anonymous):

thank you so much

OpenStudy (bahrom7893):

since the square root of the discriminant is not an integer. You're welcome :)